OpenStudy (anonymous):

which of the following is equal to 5^5/5^3

OpenStudy (anonymous):

it's not equal

OpenStudy (anonymous):

5.5.5.5.5/5.5.5=3125/125

OpenStudy (anonymous):

which of the following ... what? where are the answer choices

OpenStudy (gw2011):

5^5/5^3=5^2=25